OpenStudy (anonymous):

help with solving this? x+3/2x =5/2

OpenStudy (kamille):

5/2=2,5 \[\frac{ x+3 }{ 2x }=2,5\] \[\frac{ x+3 }{ 2x }-2,5=0\] now you need to "common denominantor". Do you know how to do it?

OpenStudy (kamille):

find*

OpenStudy (anonymous):

Can you help me go through this whole problem step by step?

OpenStudy (kamille):

Well, I can try to explain  Do you understand how I get this?